OpenStudy (anonymous):

determine the resulting temperature when 1kg of ice is mixed with 900 g of water at 50 degree celcius???

OpenStudy (anonymous):

Let the final temperature of the water be x and imagine that there is sufficient energy to melt all the ice so that x > 0 This can be confirmed or otherwise shortly. The ice requires 334000 J to melt plus 4160 * x to raise the temperature to x This energy comes from cooling the remainder of the water mass * temperature change * specific heat 9 * (50 -x ) * 4160 hence 334000 + 4160 * x = 9 * ( 50 -x ) * 4160 Now solve 334000 + 4160 x = 450 * 4160 - 9x * 4160

OpenStudy (anonymous):

hwhattt icant understand ... @TwinYang

OpenStudy (anonymous):

simplify 334000 + 4160 x = 450 * 4160 - 9x * 4160

OpenStudy (anonymous):

math error

OpenStudy (anonymous):

@TwinYang